An air-standard Ericsson cycle has an ideal regenerator. Heat is supplied at 1000oC and heat is rejected at 20oC. Pressure at the beginning of the isothermal compression is 70kPa. The heat added is 600kJ/kg. Find the compressor work, the turbine work and the cycle efficiency.
1
Expert's answer
2021-07-26T16:51:01-0400

Identify the states

Heat supplied at high temperature

"T _{3}= T _{4}=1000^{\circ} C =1273.15 K"

Heat rejected at low temperature

"T _{1}= T _{2}=80^{\circ} C =353.15 K"

Beginning of the compression:

"P _{1}=70 kPa"

Ideal regenerator:

"{ }_{2} q _{3}=-{ }_{4} q _{1} \quad \Rightarrow \quad q _{ H }={ }_{3} q _{4}=700 kJ / kg"

"m\n\n\u21d2 w _{ T }= q _{ H }= 7 0 0 k J \/ kg"



"\begin{aligned}&\eta_{ TH }=\eta_{\text {CARNOT }}=1-\frac{353.15}{1273.15}= 0 . 7 2 2 6 \\& w _{ NET }=\eta_{ TH } q _{ H }=0.7226 \times 700=505.8 kJ / kg \\& w _{ C }= q _{ L }= q _{ H }- w _{ NET }=700-505.8= 1 9 4 . 2 kJ / kg\end{aligned}"